A ball is thrown straight up into the air. The approximate height h of the ball after t seconds can be appromimately modeled by the function h=-16t^2+55t.What is the maximum height that the ball reaches?
Answer by stanbon(75887)

A ball is thrown straight up into the air. The approximate height h of the ball after t seconds can be appromimately modeled by the function
h(t) =-16t^2+55t. What is the maximum height that the ball reaches?
-----
Max occurs when t = -b/(2a) = -55/(2*-16) = 55/32
----
Height at that time = h(55/32) = 47.27 ft.
==========================
Cheers,
Stan H.